Solve equation for all value of ø if ø measure in degrees
2cos^2ø+2=5cosø

Answer by stanbon(75887)

Solve equation for all value of ø if ø measure in degrees
2cos^2ø+2=5cosø
-------------
2cos^2 - 5cos + 2 = 0
---
2cos^2 - 4cos-cos +2 = 0
---
2cos(cos-2) - (cos-2) = 0
(cos-2)(2cos-1) = 0
cos = 2 or cos = 1/2
---
cos(theta) = 1/2
----
theta = pi/3
===================
Cheers,
Stan H.
==========